# Fourier PINNs: From Strong Boundary Conditions to Adaptive Fourier Bases

### v1

@@ 1 Introduction
Physics-informed neural networks (PINNs) (Raissi et al., 2019) are innovative, mesh-free approaches for solving partial differential equations (PDEs). They offer computationally efficient alternatives to traditional mesh-based numerical methods such as finite elements and finite volumes (Reddy, 2019). The optimization of PINNs involves softly constraining neural networks (NNs) through customized loss functions designed to adhere to the governing equations of physical processes. Researchers have successfully applied PINNs in various domains—for example, they have been used to simulate the radiative transport equation (Mishra & Molinaro, 2021), which is crucial for radio frequency chip and material design  (Chen et al., 2020; Liu & Wang, 2019). Further, cardiovascular flow modeling (Kissas et al., 2020) is another application, along with various fluid mechanics problems (Cai et al., 2021), and high-speed aerodynamic flow modeling (Mao et al., 2020), among many others (Raissi et al., 2020; Chen et al., 2020; Jin et al., 2021; Sirignano & Spiliopoulos, 2018; Zhu et al., 2019; Geneva & Zabaras, 2020; Sahli Costabal et al., 2020; Sun et al., 2020; Fang & Zhan, 2019).
@@
Despite these successes, the training of PINNs remains challenging in some instances. Recent studies have analyzed common failure modes of PINNs, particularly when modeling problems that exhibit high-frequency, multi-scale, chaotic, or turbulent behaviors (Wang et al., 2022b; 2020b; 2020a; a), or when the governing PDEs are stiff (Krishnapriyan et al., 2021; Mojgani et al., 2022). Rahaman et al. (2019) attributes the slow convergence to the high-frequency components of the target solution, identifying it as a “spectrum bias” in standard NNs, while learning low-frequency information from data is straightforward. Wang et al. (2020b) confirmed that this bias is also present in the PINN setting. These challenges often arise because applying differential operators over the NN in the residual complicates the loss landscape (Krishnapriyan et al., 2021). From an optimization perspective, Wang et al. (2020a) highlighted that the imbalance in gradient magnitudes between the boundary loss and residual loss (with the latter often being much larger) causes the residual loss to dominate training, leading to a poor fit to the boundary conditions. Wang et al. (2022b) confirmed this conclusion through a neural tangent kernel (NTK) analysis of PINNs with wide networks, finding that the dominant eigenvalues of the residual kernel matrix often result in the training primarily fitting the residual loss.
@@
One class of approaches designed to mitigate the training challenges in PINNs involves setting different weights for the boundary and residual loss terms. For example, Wight & Zhao (2020) suggested incorporating a large multiplier for the boundary loss term to prevent the residual loss from dominating the training process. In contrast, Wang et al. (2020a) proposed a dynamic weighting scheme based on the gradient statistics of the loss terms, and Wang et al. (2022b) developed an adaptive weighting approach based on the eigenvalues of the NTK. Liu & Wang (2021) employed a mini-max optimization to update the loss weights via stochastic ascent, and McClenny & Braga-Neto (2020) used a multiplicative soft attention mask to dynamically re-weight the loss term for each data point and collocation point. To alleviate the spectrum bias in NNs, Tancik et al. (2020) randomly sampled a set of high frequencies from a large-variance Gaussian distribution to construct random Fourier features as input to the NN. Additionally, Wang et al. (2021) used multiple Gaussian variances to sample frequencies for Fourier features, aiming to capture multi-scale solution information within the PINN framework. While effective, the performance of this method is sensitive to the number and scales of the Gaussian variances, which are user-specified hyperparameters that are often difficult to optimize.
@@
Another strategy to improve these challenges is to modify the NN architecture to exactly satisfy the boundary conditions (BCs) (Lu et al., 2021; Lyu et al., 2020; Lagaris et al., 1998; 2000; McFall & Mahan, 2009; Berg & Nyström, 2018; Lagari et al., 2020; Mojgani et al., 2022). We refer to these approaches as “strong BC PINNs”. Despite their effectiveness, these approaches face several limitations. They are usually limited to tasks with relatively simple and well-defined physics, requiring significant craftsmanship and complex implementations even in straightforward problem settings. Consequently, these strategies are less flexible than the original PINN framework, which employs a soft constraint approach for boundary condition satisfaction. Additional complications arise for challenging physical systems governed by invariances or conservation laws, such as energy or momentum conservation, due to often poorly understood and imprecisely defined physical laws. Incorporating these laws effectively into the NN architecture makes extending strong BC PINNs to handle more complex tasks difficult. Further, designing a custom ansatz requires tailoring the NN architecture for each specific boundary condition and domain, which is often impractical for complicated domains. However, when properly designed, these techniques can achieve highly accurate solutions.
@@
  • We conduct a Fourier analysis on both strong BC PINNs compared to the standard PINN. Through the Fourier series convolution theory, we discovered that multiplying the NN by the strong boundary function significantly enhances the learning speed and accuracy of the higher frequency coefficients in the target solution. In contrast, standard PINNs struggle to capture coefficients in the high-frequency domain accurately. This analysis complements and confirms the aforementioned NTK work.
@@
The remainder of this paper is structured as follows: Section 2 provides the necessary background and notation, while in Section 4, we present our Fourier analysis of strong BC PINNs and explain the success of the strong boundary ansatz methodology. Our new Fourier PINN architecture and its training routines are described in Section 5. Section 6 details our numerical experiments and findings, including assessments of computational cost and accuracy compared to baseline methods. Finally, Section 7 discusses the results and outlines future research directions.
@@ 2.1 General Overview of Physics-Informed Neural Networks
To solve the PDE, the PINN uses a deep NN, $u_{N}({\bf x};{\bm{\theta}})$, to approximate the true solution $u({\bf x})$. For clarity in later sections, we follow the convention of Cyr et al. (2020) and define the output of the NN with width $W$ as a linear combination of a set of non-linear bases such that,
@@
where each $\bm{\psi}_{j}$ are non-linear activation functions (such as Tanh) acting on the hidden layer outputs. Each $c_{j}$ for $j=1,...,w$ and ${\bm{\theta}}^{H}$ are the weights and biases in the last layer of the NN and hidden layers, respectively. Therefore, ${\bm{\theta}}=\{\textbf{c},{\bm{\theta}}^{H}\}$ form the set of all network parameters. Then, finding the optimal network parameters ${\bm{\theta}}^{*}$ involves minimizing the following composite loss function between boundary and residual loss terms,
@@ 2.2 Strong Boundary Condition PINNs
Equation 5 can approximately enforce various types of boundary conditions, including Dirichlet, Neumann, Robin, and periodic. However, prior analysis by Wang et al. (2020a; 2022b) suggests that the instability in training PINNs likely arises from the competition between the weakly enforced boundary loss (5) and the residual loss (6) terms during optimization. This competition likely occurs because both loss terms are minimized simultaneously during training. However, the optimization algorithm tends to prioritize the minimization of the residual loss, as it typically dominates the gradient of the combined loss function. Consequently, this can lead to a scenario where the residual loss converges effectively but at the expense of the boundary loss, which remains inadequately optimized. As a result, the boundary conditions may not be satisfied, leading to poor model performance on new, unseen data or data at the domain’s boundaries.
One solution to mitigate the competition between the loss terms is to design a surrogate model that inherently satisfies the boundary conditions. This approach typically works by modifying the network architecture to satisfy the boundary conditions exactly, thus eliminating the need for a boundary enforcement loss term and further reducing the computational cost. This method of exact imposition of boundary conditions has become a standard approach in the PINN literature and is especially prevalent for Dirichlet and periodic boundary conditions (Lu et al., 2021; Yu et al., 2022; Wang et al., 2024).
@@
such that $u_{\bm{\theta}}(x)$ is the solution function, $g(x)$ is the provided boundary function, $u_{N}(x)$ is the output of the neural network, and $\phi(x)$ is a composite distance function that zeroes out on the boundaries, ensuring the boundary conditions are met without explicit penalty terms. By construction, $u_{\bm{\theta}}(a)=u_{\bm{\theta}}(b)=g(x)$, and thus the boundary condition in equation 7 is strictly and automatically satisfied. To estimate the parameters of $u_{\bm{\theta}}$, we only need to minimize the residual loss $L_{r}({\bm{\theta}})$. Most PINN surrogate model designs rely on distance functions based on the theory of R-functions (Sukumar & Srivastava, 2022), which are smooth mathematical functions that encode Boolean logic and facilitate the combination of simple shapes to form complex geometries (Rvachev & Sheiko, 1995). Some meshfree Galerkin methods have utilized the capability of R-functions to enforce boundary conditions smoothly and exactly to enhance the precision and robustness of numerical simulations for solving boundary-value problems (Shapiro & Tsukanov, 1999; Akin, 2014; Tsukanov & Posireddy, 2011).
@@
In this work, we compare the standard PINN to two strong BC PINNs using different distance functions to enforce Dirichlet boundary conditions exactly. The first distance function we consider is the polynomial distance function proposed in (Lu et al., 2021), defined as,
@@
This function satisfies the boundary conditions by zeroing out at the endpoints $x=a$ and $x=b$. While Lu et al. (2021) demonstrates promising results using this formulation, the polynomial distance function $\phi_{poly}(x)$ remains static throughout the training process. The absence of adaptability limits the model’s performance across different problems and domains, whereas introducing adaptability would allow the model to adjust to the varying complexities within the solution space.
To address this limitation, we propose an adaptive distance function to introduce flexibility in enforcing the boundary conditions, which we define as:
$\phi_{exp}(x)=(1-e^{\alpha(a-x)})(1-e^{\alpha(x-b)}),$ (10)
where $\alpha$ is a parameter that can be pre-set or optimized during training. The exponential form of equation 10 allows the function to adjust dynamically to better conform to the unique characteristics of the problem domain, thus enhancing the enforcement of boundary conditions across different scenarios and potentially leading to more accurate and efficient training outcomes.
@@ 3 Boundary Condition Pathologies in PINNs
Recent works have employed strong boundary enforcement strategies  (Lu et al., 2021; Yu et al., 2022) but offer few insights into the specific surrogate model design process. The design and implementation of the surrogate models are often ad-hoc, with no systematic approach to their architectural design. This complicates their application and limits their effectiveness across boundary conditions and physical problems, especially when dealing with complex physics and domains. Moreover, the standard PINNs frequently encounter difficulties with high-frequency and multi-scale solutions, leading to inaccurate predictions (Wang et al., 2020b). We show that while the strong BC PINN model improves solution quality for problems with higher frequency compared to the standard PINN, the solution quality nevertheless degrades with larger frequencies. In the following, we empirically demonstrate how the standard PINN solution quality degrades as the actual frequency of the solution increases in both a linear 1D Poisson problem and a non-linear 1D steady-state Allen-Cahn problem. Specifically, we consider the fabricated solution $u(x)=\sin(kx)$ for a range of frequencies $k\in[2,6,10,14,18,22,26,30,34]$ and $x\in[0,2\pi]$. Therefore, by varying $k$, we can examine the performance of the standard PINNs compared to the strong BC PINNs when the solution $u$ includes different frequency information.
@@
which gives the forcing function $f(x)=-k^{2}\sin(kx)+\sin(kx)(\sin^{2}(kx)-1)$. Both are subject to the Dirichlet condition in equation 7 where $a=0$ and $b=2\pi$. We test these problems using a standard PINN, which explicitly incorporates the weakly enforced boundary loss and residual loss terms. We compare these results to two strong BC PINNs employing different distance functions (8) and show that this method exhibits slower solution degradation with increased frequency. We used the same NN architecture for the strong BC PINN and the standard PINN, including two hidden layers, $100$ neurons per layer, and Tanh activation. Both methods used the same 10K collocation points randomly sampled from the domain. We report the averaged results for each method from five random trials.
@@
Figure 2: Relative $\ell_{2}$ error in predicting the solution $u(x)=\sin(kx)$ as a function of frequency $k$. The plots compare the performance of the standard PINN, the strong BC PINN with polynomial boundary function $\phi_{poly}$, and the strong BC PINN with exponential boundary function $\phi_{exp}$ for (left) 1D Poisson and (right) 1D steady-state Allen-Cahn equations. The shaded regions represent the error variability.
Figure 2 shows that both of the strong BC PINNs given by equation 9 and equation 10 push the surrogate model to inhibit the high-frequency components resulting in higher accuracy solutions, however, once the frequency surpasses a certain threshold, the static polynomial distance function $\phi_{poly}$ fails to provide any additional benefit. Figure 2 shows that the strong BC PINN ($\phi_{exp}$) given by equation 10 obtains higher accuracy solutions than both the standard PINN and the strong BC PINN ($\phi_{poly}$). However, it similarly hits a threshold where the benefits of the strong boundary enforcement diminish, albeit at a much higher frequency than the strong BC PINN ($\phi_{poly}$).

### v2

@@ 1 Introduction
Physics-informed neural networks (PINNs) (Raissi et al., 2019) are innovative, mesh-free approaches for solving partial differential equations (PDEs). They offer alternatives to traditional mesh-based numerical methods such as finite elements and finite volumes (Reddy, 2019). The optimization of PINNs involves softly constraining neural networks (NNs) through customized loss functions designed to adhere to the governing equations of physical processes. Researchers have successfully applied PINNs in various domains—for example, they have been used to simulate the radiative transport equation (Mishra & Molinaro, 2021), which is crucial for radio frequency chip and material design  (Chen et al., 2020; Liu & Wang, 2019). Further, cardiovascular flow modeling (Kissas et al., 2020) is another application, along with various fluid mechanics problems (Cai et al., 2021), and high-speed aerodynamic flow modeling (Mao et al., 2020), among many others (Raissi et al., 2020; Chen et al., 2020; Jin et al., 2021; Sirignano & Spiliopoulos, 2018; Zhu et al., 2019; Geneva & Zabaras, 2020; Sahli Costabal et al., 2020; Sun et al., 2020; Fang & Zhan, 2019).
@@
Despite these successes, the training of PINNs remains challenging in some instances. Recent studies have analyzed common failure modes of PINNs, particularly when modeling problems that exhibit high-frequency, multi-scale, chaotic, or turbulent behaviors (Wang et al., 2022; 2021c; 2021b; 2024b), or when the governing PDEs are stiff (Krishnapriyan et al., 2021; Mojgani et al., 2023). Rahaman et al. (2019) attributes the slow convergence to the high-frequency components of the target solution, identifying it as a “spectrum bias” in standard NNs, while learning low-frequency information from data is straightforward. Wang et al. (2021c) confirmed that this bias is also present in the PINN setting. These challenges often arise because applying differential operators over the NN in the residual complicates the loss landscape (Krishnapriyan et al., 2021). From an optimization perspective, Wang et al. (2021b) highlighted that the imbalance in gradient magnitudes between the boundary loss and residual loss (with the latter often being much larger) causes the residual loss to dominate training, leading to a poor fit to the boundary conditions. Wang et al. (2022) confirmed this conclusion through a neural tangent kernel (NTK) analysis of PINNs with wide networks, finding that the dominant eigenvalues of the residual kernel matrix often result in the training primarily fitting the residual loss.
@@
One class of approaches designed to mitigate the training challenges in PINNs involves setting different weights for the boundary and residual loss terms. For example, Wight & Zhao (2020) suggested incorporating a large multiplier for the boundary loss term to prevent the residual loss from dominating the training process. In contrast, Wang et al. (2021b) proposed a dynamic weighting scheme based on the gradient statistics of the loss terms, and Wang et al. (2022) developed an adaptive weighting approach based on the eigenvalues of the NTK. Liu & Wang (2021) employed a mini-max optimization to update the loss weights via stochastic ascent, and McClenny & Braga-Neto (2024) used a multiplicative soft attention mask to dynamically re-weight the loss term for each data point and collocation point. To alleviate the spectrum bias in NNs, Tancik et al. (2020) randomly sampled a set of high frequencies from a large-variance Gaussian distribution to construct random Fourier features as input to the NN. Additionally, Wang et al. (2021c) used multiple Gaussian variances to sample frequencies for Fourier features, aiming to capture multi-scale solution information within the PINN framework. While effective, the performance of this method is sensitive to the number and scales of the Gaussian variances, which are user-specified hyperparameters that are often difficult to optimize.
@@
Another strategy to improve these challenges is to modify the NN architecture to exactly satisfy the boundary conditions (BCs) (Lu et al., 2021; Lyu et al., 2021; Lagaris et al., 1998; 2000; McFall & Mahan, 2009; Berg & Nyström, 2018; Lagari et al., 2020; Mojgani et al., 2023). We refer to these approaches as “strong BC PINNs”. Despite their effectiveness, these approaches face several limitations. They are usually limited to tasks with relatively simple and well-defined physics, requiring significant craftsmanship and complex implementations even in straightforward problem settings. Consequently, these strategies are less flexible than the original PINN framework, which employs a soft constraint approach for boundary condition satisfaction. Additional complications arise for challenging physical systems governed by invariances or conservation laws, such as energy or momentum conservation, due to often poorly understood and imprecisely defined physical laws. Incorporating these laws effectively into the NN architecture makes extending strong BC PINNs to handle more complex tasks difficult. Further, designing a custom ansatz requires tailoring the NN architecture for each specific boundary condition and domain, which is often impractical for complicated domains. However, when properly designed, these techniques can achieve highly accurate solutions.
@@
  • We conduct a Fourier analysis on both strong BC PINNs compared to the standard PINN. Through the Fourier series convolution theory, we discovered that multiplying the NN by the strong boundary function significantly enhances the learning speed and accuracy of the higher frequency coefficients in the target solution. In contrast, standard PINNs struggle to accurately capture coefficients in the high-frequency domain. This analysis complements and confirms the aforementioned NTK work.
@@
The remainder of this paper is structured as follows: Section 2 provides the necessary background and notation, while in Section 4, we present our Fourier analysis of strong BC PINNs and explain the success of the strong boundary ansatz methodology. Our new Fourier PINN architecture and its training routines are described in Section 5. Section 6 details our numerical experiments and findings, including assessments of computational cost and accuracy compared to baseline methods. Finally, Section 7 discusses future work regarding scalability and Section 8 discusses the results and outlines specific future research directions.
@@ 2.1 General Overview of Physics-Informed Neural Networks
To solve the PDE, the PINN uses a deep NN, $u_{N}({\bf x};{\bm{\theta}})$, to approximate the true solution $u({\bf x})$. For clarity in later sections, we follow the convention of Cyr et al. (2020) and define the output of the NN with width $W$ as a linear combination of a set of nonlinear bases such that
@@
where each $\bm{\psi}_{j}$ are nonlinear activation functions (such as Tanh) acting on the hidden layer outputs. Each $c_{j}$ for $j=1,...,w$ and ${\bm{\theta}}^{H}$ are the weights and biases in the last layer of the NN and hidden layers, respectively. Therefore, ${\bm{\theta}}=\{\textbf{c},{\bm{\theta}}^{H}\}$ form the set of all network parameters. Then, finding the optimal network parameters ${\bm{\theta}}^{*}$ involves minimizing the following composite loss function between boundary and residual loss terms,
@@ 2.2 Strong Boundary Condition PINNs
Equation 5 can approximately enforce various types of boundary conditions, including Dirichlet, Neumann, Robin, and periodic. However, prior analysis by Wang et al. (2021b; 2022) suggests that the instability in training PINNs likely arises from the competition between the weakly enforced boundary loss (5) and the residual loss (6) terms during optimization. This competition likely occurs because both loss terms are minimized simultaneously during training. However, the optimization algorithm tends to prioritize the minimization of the residual loss, as it typically dominates the gradient of the combined loss function. Consequently, this can lead to a scenario where the residual loss converges effectively but at the expense of the boundary loss, which remains inadequately optimized. As a result, the boundary conditions may not be satisfied, leading to poor model performance on new, unseen data or data at the domain’s boundaries.
One solution to mitigate the competition between the loss terms is to design a surrogate model that inherently satisfies the boundary conditions. This approach typically works by modifying the network architecture to satisfy the boundary conditions exactly, thus eliminating the need for a boundary enforcement loss term and further reducing the computational cost. This method of exact imposition of boundary conditions has become a standard approach in the PINN literature and is especially prevalent for Dirichlet and periodic boundary conditions (Lu et al., 2021; Yu et al., 2022; Wang et al., 2024a).
@@
such that $u_{\bm{\theta}}(x)$ is the solution function, $g(x)$ is the provided boundary function, $u_{N}(x)$ is the output of the neural network, and $\phi(x)$ is a composite distance function that zeroes out on the boundaries, ensuring the boundary conditions are met without explicit penalty terms. By construction, $u_{\bm{\theta}}(a)=u_{\bm{\theta}}(b)=g(x)$, and thus the boundary condition in Equation 7 is strictly and automatically satisfied. To estimate the parameters of $u_{\bm{\theta}}$, we only need to minimize the residual loss $L_{r}({\bm{\theta}})$. Most PINN surrogate model designs rely on distance functions based on the theory of R-functions (Sukumar & Srivastava, 2022), which are smooth mathematical functions that encode Boolean logic and facilitate the combination of simple shapes to form complex geometries (Rvachev & Sheiko, 1995). Some meshfree Galerkin methods have utilized the capability of R-functions to enforce boundary conditions smoothly and exactly to enhance the precision and robustness of numerical simulations for solving boundary-value problems (Shapiro & Tsukanov, 1999; Akin, 1994; Tsukanov & Posireddy, 2011).
@@
In this work, we compare the standard PINN to two strong BC PINNs using different distance functions to enforce Dirichlet boundary conditions exactly. The first distance function we consider is the polynomial distance function proposed by (Lu et al., 2021), defined as,
@@
This function satisfies the Dirichlet boundary conditions by zeroing out at the $x=a$ and $x=b$ endpoints. While Lu et al. (2021) demonstrated promising results using this formulation, the polynomial distance function $\phi_{poly}(x)$ remains static throughout the training process. The lack of adaptability limits the model’s ability to handle different problem domains and boundary conditions, which can hinder performance when faced with varying complexities within the solution space.
To address this limitation, we propose an adaptive distance function that introduces flexibility in enforcing boundary conditions, which we define as:
$\phi_{exp}(x)=(1-e^{\alpha_{a}(a-x)})(1-e^{\alpha_{b}(x-b)}),$ (10)
where $\alpha_{a}$ and $\alpha_{b}$ are parameters that can be pre-set or optimized during training. These parameters allow the function to adjust to the characteristics of the problem domain dynamically, ensuring robust enforcement of boundary conditions across different scenarios. As we demonstrate in subsequent sections, this flexibility improves accuracy and enhances the efficiency of the training process by enabling the network to adapt to the complexity of the solution space. Allowing independent values for $\alpha_{a}$ and $\alpha_{b}$ enables additional flexibility and allows the function to adapt asymmetrically to different boundary conditions. Although the proposed adaptive distance function in Equation 10 may appear incremental, its primary contribution lies in demonstrating a scalable and flexible framework for enforcing strong boundary conditions. This method significantly improves over traditional PINNs and non-adaptive distance functions, particularly for problems that require the exact imposition of Dirichlet boundary conditions. Additionally, in later sections, we present a thorough theoretical analysis that elucidates the mechanisms driving the improvements in enforcing strong boundary conditions, further highlighting the advantages of the proposed method.
The general form of the adaptive distance function for a $d$-dimensional rectilinear domain $\Omega\subset\mathbb{R}^{d}$, with spatial coordinates $\mathbf{x}=(x_{1},x_{2},...,x_{d})$, is:
$\phi_{\text{exp}}(\mathbf{x})=\prod_{i=1}^{d}\left(1-e^{\alpha_{a_{i}}(a_{i}-x_{i})}\right)\left(1-e^{\alpha_{b_{i}}(x_{i}-b_{i})}\right),$ (11)
where $a_{i}$, $b_{i}$ represent the boundaries in the $i$-th dimension, and $\alpha_{a_{i}},\alpha_{b_{i}}$ are the corresponding steepness parameters. This form ensures that the boundary conditions are enforced independently in each dimension, making the approach easily scalable to higher dimensions in rectilinear domains.
One potential strategy for adaptation for irregular domains is using Fourier extension methods, which allow non-periodic functions defined on arbitrary domains to be extended into periodic ones. Doing so can enforce the boundary conditions in a more structured domain, allowing the network to handle irregularities better. Another promising direction is leveraging custom network layers that map irregular domains to rectilinear spaces, as demonstrated by methods such as the GOFNO network Liu et al. (2023). Extending this framework to more complex boundary conditions, such as Neumann or Robin conditions, is an important future direction. The adaptive distance function can be modified for Neumann boundary conditions to have non-zero slopes at the boundaries, ensuring correctly enforced derivatives. This would require careful and non-trivial adjustments to the distance function, but the general framework remains applicable.
@@ 3 Boundary Condition Pathologies in PINNs
Recent works have employed strong boundary enforcement strategies  (Lu et al., 2021; Yu et al., 2022) but offer few insights into the specific surrogate model design process. The design and implementation of the surrogate models are often ad-hoc, with no systematic approach to their architectural design. This complicates their application and limits their effectiveness across boundary conditions and physical problems, especially when dealing with complex physics and domains. Moreover, the standard PINNs frequently encounter difficulties with high-frequency and multi-scale solutions, leading to inaccurate predictions (Wang et al., 2021c). We show that while the strong BC PINN model improves solution quality for problems with higher frequency compared to the standard PINN, the solution quality nevertheless degrades with larger frequencies. In the following, we empirically demonstrate how the standard PINN solution quality degrades as the actual frequency of the solution increases in both a linear 1D Poisson problem and a non-linear 1D steady-state Allen-Cahn problem. Specifically, we consider the fabricated solution $u(x)=\sin(kx)$ for a range of frequencies $k\in[2,6,10,14,18,22,26,30,34]$ and $x\in[0,2\pi]$. Therefore, by varying $k$, we can examine the performance of the standard PINNs compared to the strong BC PINNs when the solution $u$ includes different frequency information.
@@
which gives the forcing function $f(x)=-k^{2}\sin(kx)+\sin(kx)(\sin^{2}(kx)-1)$. Both are subject to the Dirichlet condition in Equation 7 where $a=0$ and $b=2\pi$.
@@
We test these problems using a standard PINN, which explicitly incorporates the weakly enforced boundary loss and residual loss terms. We compare these results to two strong BC PINNs employing different distance functions (8) and show that this method exhibits slower solution degradation with increased frequency. Specifically, we tested two neural network architectures, each with $100$ neurons per layer and depths of two and four, and assessed model performance using the relative $\ell_{2}$ error, defined as
$\frac{\|y-\hat{y}\|_{2}}{\|y\|_{2}}=\frac{\sqrt{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}}{\sqrt{\sum_{i=1}^{n}y_{i}^{2}}},$
where $\hat{y}$ is the predicted solution and $y$ is the true solution. All networks used the Tanh activation and are trained using $10$K equally spaced collocation points sampled from the domain. Relative $\ell_{2}$ errors are reported on a separate testing set of $20$K points. All experiments used 32-bit floating-point precision (float32) to ensure computational efficiency. Each model was trained for $100$K iterations using the Adam optimizer (Kingma & Ba, 2015) with an initial learning rate of $10^{-3}$, decaying exponentially by $0.9$ every $1000$ iterations, followed by L-BFGS optimization until convergence (tolerance $10^{-9}$). For the strong BC PINN with $\phi_{\text{exp}}$, we initialized $\alpha$ to $0.5$ and jointly optimized it with all other parameters. Results are averaged over five random trials and were conducted on a GeForce RTX 3090 GPU with CUDA version 12.3, running on Ubuntu 20.04.6 LTS. Additionally, we implemented a second-order accurate Finite Difference Method (FDM) as a baseline for solving the 1D Poisson and steady-state Allen-Cahn equations using 1000 discretization points and central differences. Table 3 in the Appendix provides detailed hyperparameters and experimental design information.
@@
Figure 2: Relative $\ell_{2}$ error in predicting the solution $u(x)=\sin(kx)$ as a function of frequency $k$. The plots compare the performance of the standard PINN, the strong BC PINN with polynomial boundary function $\phi_{poly}$, the strong BC PINN with exponential boundary function $\phi_{exp}$, and FDM for (left) 1D Poisson and (right) 1D steady-state Allen-Cahn equations for NN with $2$ hidden layers (d-2) and $4$ hidden layers (d-4) each with $100$ neurons. The shaded regions represent the error variability.
Figure 2 demonstrates that both strong BC PINNs, using $\phi_{\text{poly}}$ and $\phi_{\text{exp}}$ boundary functions, effectively suppress high-frequency components, leading to higher accuracy solutions compared to the standard PINN. The deeper architecture (d=4) consistently reduces errors across most frequencies for both boundary functions, but the static polynomial distance function $\phi_{\text{poly}}$ shows diminishing benefits as the frequency increases. In contrast, the strong BC PINN with $\phi_{\text{exp}}$ maintains superior accuracy, outperforming both the standard PINN and the strong BC PINN with $\phi_{\text{poly}}$ across a broader frequency range. Although the advantages of strong boundary enforcement diminish at very high frequencies, $\phi_{\text{exp}}$ consistently retains its edge over $\phi_{\text{poly}}$.
Additionally, the Finite Difference Method (FDM) achieves the lowest relative $\ell_{2}$ errors across all frequencies in both the 1D Poisson and 1D Allen-Cahn problems, serving as a benchmark for evaluating neural network approaches. The FDM demonstrates exceptional performance, particularly at high frequencies, highlighting the limitations of PINN-based approaches in capturing rapid oscillations. However, the strong BC PINNs with $\phi_{\text{exp}}$ narrow the gap with FDM at lower and mid-range frequencies, showcasing their robustness compared to standard PINNs.
While the FDM is a reliable and efficient approach for solving partial differential equations, it has notable limitations that drive the need for advancements in Physics-Informed Neural Networks (PINNs). FDM relies on structured grids, making it challenging to apply to problems with complex geometries, and it struggles with scalability in high-dimensional problems due to exponential grid growth. Unlike FDM, PINNs can also integrate noisy or partial data into their models, making them more flexible for real-world applications. Furthermore, advancements in machine learning architectures and hardware accelerators have significantly improved the practicality and efficiency of PINNs. By addressing current challenges, such as resolving high-frequency solutions and improving convergence, PINNs have the potential to provide a scalable, versatile alternative to traditional numerical methods like FDM.
Table 1: Training times (in seconds) for various network configurations across 1D Poisson and 1D Allen-Cahn problems. The values in parentheses represent the standard deviations of training times.
Network 1D Poisson (sec) 1D Allen-Cahn (sec)
Standard PINN (d=2) $202.07\,(2.52)$ $103.11\,(0.10)$
Standard PINN (d=4) $319.09\,(1.61)$ $200.14\,(91.20)$
Strong BC PINN $\phi_{\text{poly}}$ (d=2) $204.28\,(1.95)$ $213.41\,(78.60)$
Strong BC PINN $\phi_{\text{poly}}$ (d=4) $311.83\,(10.83)$ $122.91\,(15.90)$
Strong BC PINN $\phi_{\text{exp}}$ (d=2) $299.56\,(2.18)$ $403.60\,(260.00)$
Strong BC PINN $\phi_{\text{exp}}$ (d=4) $401.63\,(1.98)$ $472.59\,(56.00)$
FDM $<1$ $<1$